# Mind the Privacy Unit! User-Level Differential Privacy for Language Model Fine-Tuning

### v2

@@ 8 Conclusions
While our work provides valuable insights, we acknowledge a few limitations and suggest directions for future research. First, in our implementation of User-wise DP-SGD, we fixed the value of $k$ (the number of gradients to sample per user) for simplicity. Exploring more dynamic choices of $k$ may yield better utility while preserving privacy guarantees. Second, since User-wise DP-SGD can perform data selection on the fly, it will be interesting to investigate how other data selection methods, such as those based on gradient information or active learning, might impact the utility. Finally, while we focus on language generation, user-level DP has broader applications motivating analysis in domains like recommendation systems and structured prediction.
Appendix A Algorithm
We describe the algorithm in Asi & Liu (2024) in our notation. They use full-batch size.
@@ B.2 Training details
Privacy Accounting.
We use the privacy accounting algorithm based on privacy loss distributions (PLD) as implemented in the open-source dp_accounting library (Google., 2020). This privacy accounting assumes that the batches are constructed using Poisson subsampling. Even though our implementation involves constructing batches by going through users/records in a random order, we use the accounting assuming Poisson subsampling as is common practice in the privacy accounting of DP-SGD.

### v3

@@ 8 Conclusions
While our work provides valuable insights, we acknowledge a few limitations and suggest directions for future research. First, in our implementation of User-wise DP-SGD, we fixed the value of $k$ (the number of gradients to sample per user) for simplicity. Exploring more dynamic choices of $k$ for different users may yield better utility while preserving privacy guarantees. Second, since User-wise DP-SGD can perform data selection on the fly, it will be interesting to investigate how other data selection methods, such as those based on gradient information or active learning, might impact the utility. Finally, while we focus on language generation, user-level DP has broader applications motivating analysis in domains like recommendation systems and structured prediction.
Appendix A User-Level Algorithm of Asi & Liu (2024)
We describe the algorithm in Asi & Liu (2024) using our notation for completeness. They assume each user owns the same amount of record and use the full batch size (use all dataset in each step). We modify the algorithm by drawing $k_{i}^{\prime}$ records for the $i$th user in the following Pseudocode.
@@ B.2 Model details
B.3 Privacy Accounting
We perform privacy accounting assuming that the batches are constructed using Poisson subsampling, where each batch is independently sampled by including each record / user independently with a certain probability, even though our implementation involves constructing batches by going through users/records in a random order. We note that this assumption has been a common practice in the privacy accounting of DP-SGD, starting with the work of Abadi et al. (2016). Recent work by Chua et al. (2024) points out that there can be substantial gaps between the privacy guarantees of DP-SGD when using Poisson subsampling vs. shuffling; however, since we make the Poisson subsampling assumption for both Group Privacy and User-wise DP-SGD, we consider our evaluation to still be informative in a relative sense even in the presence of this gap.
The privacy analysis of DP-SGD using Poisson subsampling with $M$ records and expected batch size $B$, run for $T$ steps is performed by analyzing the $T$-fold composition of the Poisson subsampled Gaussian mechanism with subsampling probability of $B/M$. It is possible to numerically compute the smallest $\sigma$, up to some tolerance, such that this mechanism satisfies $({\varepsilon},\delta)$-DP using algorithms based on privacy loss distributions (PLD) (Koskela et al., 2021; Gopi et al., 2021; Doroshenko et al., 2022), using various open source libraries Microsoft. (2021); Prediger & Koskela (2020). In particular, we use the implementation in the dp_accounting library (Google’s DP Library., 2020). We use $\textsc{PrivAccounting}({\varepsilon},\delta,M,B,T)$ to refer to this smallest $\sigma$.
PrivAccounting for Algorithm 2
Under the assumption of Poisson subsampling of users, the same privacy analysis is applicable with “number of records” $M$ replaced by “number of users” $N$ and “batch size” $B$ replaced by “user batch size” $n$, and thus, the optimal noise parameter is given as $\textsc{PrivAccounting}({\varepsilon},\delta,N,n,T)$.
GroupPrivAccounting for Algorithm 1.
Ganesh (2024) showed that the privacy analysis of a single step of DP-SGD under addition/removal of $k$ records under Poisson subsampling is given by considering the Gaussian mechanism with sensitivity that is randomly drawn from the Binomial distribution $\mathrm{Bin}(k,B/M)$ referred to as the “Mixture of Gaussians (MoG)” mechanism. The privacy analysis of the entire DP-SGD can thus be performed by analyzing the $T$-fold composition of this MoG mechanism, which can be done in a tight manner up to numerical accuracy using privacy loss distributions. We use the implementation available in the dp_accounting library.